UT 20170119 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, DThompson (Tuson)
Observer AO: GTayler, DMiller
Telescope Operator: SAllanson

Plan for the night

- Waiting for the weather to improve, if on sky:

- BB AFC-tests

If closed:

- Tests by Dave FM1/4

- AFC-ima LUCI1

- Handover to MPE-observers at 7.30


- Severe AO communication issues, no AO-com or AO-science


All times are UT
Files are luci1.20170119.NNNN.fits

00:35 Sunset, Weather: overcast (fog) but starts to clear up, T=-7 deg, Hum=96%, Wind=low wind

01:00 No ice on dome, Steve could open 3/4 vent doors, ventilation still off due to humidity. If humidity drops
we can open, but expect lots of dome seeing (Tmirror = 2deg)

01:15 Dave starts FM1/4 tests

02:30 Complelety clear outside but humidity still high

04:15 Believe it or not, humidity just below 90%, we give it a try - Steve opens

04:30 Conditions not too bad, seeing about 1.2" on guider, DIMM reports 1.2-1.5"

Only 3hrs left, so we go for quick AO-tests for the AO-group where we may
benefit out of it in March, mostly done by Doug

04:40 Test for pupil shift (wobble) during large offsets within patrol field (loop than opens).
This is a new feature which apparently came up a few weeks ago the first time
=> still present !


05:40 Preset to Palomar 1 (Palomar1_AFC_Start1.xml), AO-ref about 15.3 (13.0 at HST-catalog),
RunAO failed on the AO supervisor, unknown exception from arbitrator,
resent preset, unknown exception from arbitrator again, seems to be
communication problem, all processes hang, low-level communication problem between AOS and
AO supervisor, both ADSec and WFS have been powered off and restartet

06:25 Resent preset, unknown exception from arbitrator again,

Communication problems severe, stopped LUCI1-AO

06:50 Handover to MPE-team for LBTB-science
Topic revision: r4 - 19 Jan 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback